OpenStudy (anonymous):

What is 1/x - 1/y over 1/x^2 - 1/ y^2 equal to?

OpenStudy (anonymous):

multiply top and bottom by \[x^2y^2\] to get the answer in the fewest steps

OpenStudy (anonymous):

or factor the denominator as the difference of two squares and cancel with the numerator.

OpenStudy (anonymous):

depending on how fancy you want to be

OpenStudy (anonymous):

or note that 1/x^2 - 1/y^2 = (1/x+1/y)(1/x-1/y)

OpenStudy (anonymous):

\[\frac{\frac{1}{x}-\frac{1}{y}}{\frac{1}{x^2}-\frac{1}{y^2}}\] \[=\frac{\frac{1}{x}-\frac{1}{y}}{(\frac{1}{x}-\frac{1}{y})(\frac{1}{x}+\frac{1}{y})}\] \[=\frac{1}{(\frac{1}{x}+\frac{1}{y})}\]

OpenStudy (anonymous):

now multiply top and bottom by \[xy\] to get \[\frac{xy}{y+x}\]

OpenStudy (anonymous):

Thank you